OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). y = -4x + 1 y = -1/2 x + 9 y = 4x + 1 y = -2x + 9

OpenStudy (anonymous):

0/-1 as m?????

OpenStudy (anonymous):

first find slope: \[m=\left({y_2-y_1}\over{x_2-x_1}\right)~~\implies~~m=\left({-1-1}\over{5-4}\right)~~\implies~~m=\left({-2}\over{1}\right)~~\implies~~m=-2\] now that we have slope plug the it into the formula \(y-y_1=m(x-x_1)\) along with either of the two provided set of points for \(x_1\) and \(y_1\) as follows:\[y-1=-2(x-4)~~\implies~~y-1=-2x+8~~\implies~~y=-2x+9\] \[\text{hope this is helpful! :)}\]

OpenStudy (anonymous):

Y = Mx + c; where M = (y2-y1)/(x2-x1) Slope is th change in x axis divided by change in x axis and x1, x2 y1 y2 are the points on the line here M= (-1-1)/(5-4) =-2/1=-2 now we need to find out the value of c. which is th intercept ... a point at which line crosses the axis substituting slope values and any point on the line {here i selected (4,1)}we get 1=-2x4+c simplifying we get c = 9...... this implies that equation of the line is Y=-2x+9 So your 4th option is correct

OpenStudy (anonymous):

thanks both of you ! i dont want to be unfair bcs you both did a lot of work d i cant give a best answer sorry

OpenStudy (anonymous):

lol no problem...anytime! :)